### v1

@@ IV-B Score calibration
Like all scores, equidensity scores can be calibrated to a desired range of values, for example by raising to an exponent (for scores that range from 0 and 1). Importantly, note that for a given scoring statistic, calibration cannot change which of two fits has the higher value. For example, no calibration procedure can make the correlation score for Fig. 2e larger than for Fig. 2a (Table I). This is a structural problem with the correlation score (and as we have shown, it affects the equipoint scores as well). This means that even though all correlation scores can be brought arbitrarily low in hopes of resolving the absolute grade inflation they exhibit (for example by raising the correlation score to a sufficiently large exponent), structurally, the correlation score will still always exhibit relative grade inflation, in which low-quality fits will perform as well as or better than high-quality ones. Thus, calibration cannot fix grade inflation. The goal of the present work has been to investigate scores’ structural properties. Calibration of the Eden score or other scores, if desired, is left for future work.
IV-C Variance and confidence
@@
Consideration of variance raises the question of whether the Eden score should always be calculated with five annuli. We propose the choice should depend on the total number of datapoints, such that sufficiently many points lie in each annulus to provide for acceptable sampling variation and acceptable confidence intervals. Because sampling error scales as the root of the number of points, we suggest one approach is to choose the number of annuli so that there are roughly 30 points per annulus; more points per annulus will result in narrower confidence intervals, but more annuli will result in a higher-resolution comparison of the two distributions. This is essentially a precision-accuracy trade-off that is the choice of the investigator. (In this study we used five annuli throughout, to control for any effects this choice might have.) For general applications beyond synthetic data in which the number of points is non-limiting—for example, when the two distributions are continuous and defined at every point—it could be interesting to define a continuum limit for $n_{\text{annuli}}\rightarrow\infty$. This is left for future work.
@@ IV-E Limitations and conclusions
The primary limitations of this work are its focus on examples of two-dimensional distributions and a small number of scores; it does not investigate higher dimensions nor attempt to systematically discover and/or evaluate all possible equipoint or equidensity scores. We note that these scores do apply in higher dimensions; however as dimensionality rises, there is risk that the curse of dimensionality will affect equidensity scores as it does, for example, the Jaccard score, leading to artifactually lower scores in higher dimensions than one might consider reasonable (in the case of the Jaccard score, because the union grows much faster than the intersection as dimensionality increases). (How a one-dimensional version of the Eden score would compare to the KS statistic or a one-dimensional KL score, is an interesting question.) Studies of higher dimensions might be considered useful, since non-trivial real datasets are often high-dimensional (e.g., many columns), but our approach of using humans as the gold standard would be complicated by the inability to easily visualize higher dimensions and potentially important losses if dimensionality reduction is used (e.g. PCA, tSNE, UMAP) [10, 9, 8]. Fortunately, lower-order relationships often carry a large amount of information in complex systems [15, 16]. For this reason, we expect Eden and other equidensity scores to be useful additions to the generative-modeling toolkit.

### v2

@@ IV-B Score calibration
Like all scores, equidensity scores can be calibrated to a desired range of values, for example by raising to an exponent (for scores that range from 0 and 1). Importantly, note that for a given scoring statistic, calibration cannot change which of two fits has the higher value. For example, no calibration procedure can make the correlation score for Fig. 2e larger than for Fig. 2a (Table I). This is a structural problem with the correlation score; as we have shown, it affects the equipoint scores as well. This means that even though all correlation scores can be brought arbitrarily low in hopes of resolving the absolute grade inflation they exhibit, for example by raising the correlation score to a sufficiently large exponent, structurally the correlation score will still always exhibit relative grade inflation, in which low-quality fits will perform as well as or better than high-quality ones. Thus, calibration cannot fix grade inflation. The goal of the present work has been to investigate scores’ structural properties. Calibration of the Eden score or other scores, if desired, is left for future work.
IV-C Sampling error and confidence
@@
Consideration of variance raises the question of whether the Eden score should always be calculated with five annuli. We propose the choice should depend on the total number of datapoints, such that sufficiently many points lie in each annulus to provide for acceptable sampling variation and acceptable confidence intervals. Because sampling error scales as the root of the number of points, we suggest one approach is to choose the number of annuli so that there are at least roughly 30 points per annulus; more points per annulus will result in narrower confidence intervals, but more annuli will result in a higher-resolution comparison of the two distributions. We believe this is essentially a precision-accuracy trade-off that is the choice of the investigator. (In this study we used five annuli throughout, to control for any effects variation in this choice might have.) For general applications beyond synthetic data in which the number of points is non-limiting—for example, when the two distributions are continuous and defined at every point—it could be interesting to define a continuum limit for $n_{\text{annuli}}\rightarrow\infty$. This is left for future work.
@@ IV-E Limitations and conclusions
The primary limitations of this work are its focus on examples of two-dimensional distributions and a small number of scoring methods; it does not investigate higher dimensions nor attempt to systematically discover and/or evaluate all possible equipoint or equidensity scores. We note that these scores do apply in higher dimensions; however as dimensionality rises, there is risk that the curse of dimensionality will affect equidensity scores as it does, for example, the Jaccard score, leading to artifactually lower scores in higher dimensions than one might consider reasonable (in the case of the Jaccard score, because the union grows much faster than the intersection as dimensionality increases). How a one-dimensional version of the Eden score would compare to the KS statistic or a one-dimensional KL score, is an interesting question. Studies of higher dimensions might be considered useful, since non-trivial real datasets are often high-dimensional (e.g., many columns), but using human raters as the gold standard would be complicated by the inability to easily visualize higher dimensions and by potentially important losses if dimensionality reduction is used (e.g. PCA, tSNE, UMAP) [10, 9, 8]. Fortunately, lower-order relationships often carry a large amount of information about the configuration of complex systems [15, 16]. For this reason, we expect Eden and other equidensity scores to be useful additions to the generative-modeling toolkit.